OpenStudy (anonymous):

\[Derivation.of ==> e ^{5x ^{4}}\]

OpenStudy (anonymous):

\[= e^{5x^4} \times \frac{d}{dx}(5x^4)\] Try to solve this..

OpenStudy (anonymous):

\[\frac{d}{dx} (5x^4) = 4 \times 5x^3\]

OpenStudy (cwrw238):

use the chain rule let u = 5x^4 y = e^u ----> dy/du = e^u = e^(5x^4) du /dx = 20x^3 dy/dx = e^(5x^4) * 20x^3

OpenStudy (anonymous):

Remember: \[\huge \color{green}{\frac{d}{dx} (e^{f(x)}) = e^{f(x)} \times \frac{d}{dx}(f(x))}\]